The parallel combination of five cells, each with electromotive force of 1.5 V and internal resistance of 0.5 Ω, is connected to an external resistance of 5.0 Ω. Find the (a) terminal voltage of the battery, and (b) terminal voltage of a cell.




Given: n=5



E= 1.5 V



r= 0.5 Ω



R= 5.0 Ω

1
Expert's answer
2022-04-21T13:00:50-0400

(a)


"R_{tot}=\frac{r}{5}+R=\frac{0.5}{5}+5.0=5.1\:\Omega"

(b)


"{\cal E}=1.5\:\rm V"

(c)


"I=\frac{\cal E}{R_{tot}}=\frac{1.5}{5.1}=0.29\:\rm A"